OpenStudy (anonymous):

A trip to Portland, Oregon from Boston will take 7 3/4 hours. Assuming we are 2/3 of the way there, how much longer in hours will the trip take? (Points : 4)

OpenStudy (anonymous):

so since we are 2/3 of the way there, you have 1/3 of the journey to travel. Thus the trip will take another 7 3/4 * 1/3 = 31/12 hours or 155 minutes

OpenStudy (anonymous):

7 3/4 * 2/3 = 31/6